3 5/6 + 5 2/3 and reduce to lowest form
Answer #1

3 5/6 = (6*3 + 5)/6 = 23/6

5 2/3 = (5*3+2)/3 = 17/3

23/6 + 17/3 = (23+34)/6

= 57/6

Cancelling by 3, 19/2 = 9 1/2